# Self-bound quark stars with a first-order two-to-three flavor phase transition

### v1

@@ IV NUMERICAL RESULTS FOR THE EOS
Table 2: First-order $ud\!\to\!uds$ transition parameters for the flavor-dependent QMDD EOS. For each $(a,C,\kappa)$ we list the transition pressure $p_{\mathrm{tr}}$, the energy densities just below/above the transition ($\epsilon_{-}$, $\epsilon_{+}$), and the jump $\Delta\epsilon\equiv\epsilon_{+}-\epsilon_{-}$. The last column reports the Seidov diagnostic $\mathit{q}\equiv(\epsilon_{+}/\epsilon_{-})\big/\bigl[\tfrac{3}{2}\bigl(1+p_{\mathrm{tr}}/\epsilon_{-}\bigr)\bigr]$; values $\mathit{q}<1$ indicate that the Seidov instability criterion is not met.
$a$ $C$ $\kappa$ $p_{\mathrm{tr}}$ (MeV fm-3) $\epsilon_{+}$ (MeV fm-3) $\epsilon_{-}$ (MeV fm-3) $\Delta\epsilon$ (MeV fm-3) $\mathit{q}$
@@
Seidov [Seidov1971] showed that, within general relativity, the onset of a first-order phase transition renders a spherically symmetric star unstable when the energy-density jump exceeds the critical value
@@
where $p_{\mathrm{tr}}$ is the transition pressure and $\epsilon_{-}$ ($\epsilon_{+}$) is the energy density just below (above) the transition. Table 2 lists $p_{\mathrm{tr}}$, $\epsilon_{-}$, and $\epsilon_{+}$ for all parameter sets considered here and, in the last column, the diagnostic ratio
$\mathit{q}\equiv\frac{\epsilon_{+}/\epsilon_{-}}{\tfrac{3}{2}\bigl(1+p_{\mathrm{tr}}/\epsilon_{-}\bigr)}\,.$ (7)
All values lie in the range 0.850.850.970.97, i.e., below unity, so Seidov’s instability criterion is not met. Hybrid configurations featuring a two- to three-flavor transition are therefore expected to remain dynamically stable; we confirm this in Sec. V by integrating the stellar structure equations and exhibiting stable branches in the mass–radius relation.
Table 2 further shows that, for fixed $(a,C)$, $p_{\mathrm{tr}}$, $\epsilon_{-}$, and $\epsilon_{+}$ vary appreciably with $\kappa$, whereas the combination in Eq. (6) is essentially insensitive to $\kappa$.
@@ V Global properties of Hybrid self-bound quark stars
Figure 5: (a) Schematic mass–radius diagram of self-bound stars. The magenta curve shows self-bound strange-quark-matter configurations, the blue curve shows self-bound udud stars, and the bicolor (magenta–blue) curve displays the sequence of self-bound hybrid stars analyzed in this work. (b) Representation of the internal composition of the models on the MMRR curves in panel (a), indicating regions of udsuds matter (magenta) and udud matter (blue).
@@
Figure 6: Mass–radius relations for quark-star sequences computed with the EOS sets of Table 1. Line styles encode the excluded-volume strength $\kappa$ (solid, dashed, dotted), following the convention of Fig. 2. Colors indicate composition: magenta—pure udsuds (SQM); blue—pure udud; bicolor—self-bound hybrids with a udsuds core surrounded by a udud mantle. (a) Parameters in the SQM domain yield purely three-flavor sequences. (b)–(d) EOSs featuring a first-order $ud\!\to\!uds$ transition placed at low, intermediate, and high densities, respectively; the kink marks $p_{c}=p_{\mathrm{tr}}$. Increasing $\kappa$ stiffens the EOS and raises $M_{\max}$. Shaded contours denote observational constraints from GW170817 [LIGOScientific:2017vwq, LIGOScientific:2018cki, LIGOScientific:2018hze], PSR J0030+0451 [Riley2019, Miller2019], HESS J1731-347 [Doroshenko2022], and PSR J0740+6620 [Riley2021, Miller2021].
@@
Figure 6 confronts the mass–radius sequences with current astrophysical constraints. A clear dependence on the excluded-volume strength is evident. Models with $\kappa=0$ (solid) are too soft: their MMRR curves fail to reach $2\,M_{\odot}$, underscoring the need for repulsive interactions in the quark sector. At the other extreme, $\kappa=0.6$ (dotted) produces very stiff EOSs. Although these sequences comfortably attain large maximum masses, they generally overpredict radii; in panels 6(b)–6(d) the curves lie on or above the upper edge of the PSR J0030+0451 and PSR J0740+6620 constraints. An exception occurs in panel 6(a), where the pure SQM models with $\kappa=0.6$ are compatible with all bands, a marked improvement over $\kappa=0$ and $0.3$. For the self-bound hybrid cases, the intermediate choice $\kappa=0.3$ (dashed) provides the best overall agreement, balancing the $2\,M_{\odot}$ requirement against radius constraints. Across panels 6(b)–6(d), the predicted $M_{\max}$ varies with the transition density: it is lower for both low and high transition densities and peaks for the intermediate case in panel 6(c).
Figure 7: Dimensionless tidal deformability $\Lambda$ as a function of gravitational mass $M$ for the same EOS sets used in Fig. 6. Line styles encode the excluded-volume strength $\kappa$ (solid, dashed, dotted), following the convention of Fig. 2. (a) Pure strange–quark–star sequences; (b)–(d) self-bound hybrid-star sequences with the $ud\!\to\!uds$ transition placed at low, intermediate, and high densities, respectively. The point with error bar indicates the GW170817 constraint at $1.4\,M_{\odot}$ [LIGOScientific:2017vwq, LIGOScientific:2018cki, LIGOScientific:2018hze].
@@
Figure 7 displays the dimensionless tidal deformability $\Lambda$, which measures the induced quadrupole per unit external tidal field and provides a direct handle to confront the EOS with gravitational-wave observations (e.g., GW170817). We compute $\Lambda(M)$ for the same four EOS sets as in the preceding figures using the formalism summarized in Appendix B.2. The point with error bar marks the GW170817 constraint at $1.4\,M_{\odot}$. Two robust trends emerge. First, for any given EOS, $\Lambda$ decreases monotonically with $M$ as stars become more compact. Second, at fixed $M$, larger excluded volume (larger $\kappa$) yields larger $\Lambda$, reflecting EOS stiffening and the associated increase in radius. Quantitatively, at fixed mass one may write $\Lambda\simeq\tfrac{2}{3}\,k_{2}\,(R/M)^{5}$, so the strong compactness dependence implies that even modest radius changes (few percent) can amplify $\Lambda$ by tens of percent.
The observational implications differ for the two classes of models. For pure strange-quark stars [Fig. 7(a)], predicted deformabilities are generally too small to satisfy GW170817; only the stiffest case ($\kappa=0.6$) attains marginal consistency at $1.4\,M_{\odot}$. Self-bound hybrid stars [Figs.7(b)–7(d)] are, in general, consistent with the GW170817 constraint over a broad range of parameters. Two notable exceptions arise: (i) for high transition density combined with a large excluded volume ($\kappa=0.6$), the predicted $\Lambda$ at $1.4,M_{\odot}$ is too large and lies outside the allowed GW170817 interval [Fig.7(b)]; and (ii) for low transition density with vanishing excluded volume ($\kappa=0.0$), the predicted $\Lambda$ at $1.4,M_{\odot}$ is too small and likewise falls outside the allowed range [Fig. 7(d)].
@@ VI Universal relations
Figure 10: (a) Dimensionless moment of inertia $I/M_{G}^{3}$ vs. compactness $M_{G}/R$ for sequences with different values of $a$. Only EOSs with a maximum mass $2\leq M_{\max}/M_{\odot}\leq 3$ are included. The dashed lines are the global fits for strange quark stars [Lugones2025universality] and hadronic stars [Breu:2016ufb]. (b) Relative error, $\bigl|1-(I/M_{G}^{3})_{\rm num}/(I/M_{G}^{3})_{\rm fit}\bigr|$, with respect to the strange quark star fit.
Figure 11: Tidal deformability $\Lambda$ as a function of the gravitational compactness $\mathscr{C}\equiv M_{G}/R$ for stellar sequences labeled by the parameter $a$. Panel (a): solid curves show the models; the cyan dashed line is the strange–quark–star fit [Lugones2025universality] and the crimson dashed line is an empirical hadronic–star fit  [Breu:2016ufb]. Panel (b): relative error with respect to the strange–quark–star fit, $|1-\Lambda/\Lambda_{\mathrm{fit}}^{\mathrm{SQS}}|$.
@@ VI.1 Dimensionless moment of inertia
Figure 10 shows the dimensionless moment of inertia, $\bar{I}\equiv I/M^{3}$, as a function of the compactness $\mathscr{C}\equiv M/R$. Despite the strong $\kappa$-dependence seen in individual MRM\!-\!R and MIM\!-\!I relations, its imprint is nearly washed out in the $\bar{I}$$\mathscr{C}$ plane: for fixed $(a,C)$, the three $\kappa$ tracks collapse to an excellent approximation onto a single curve. Moreover, the global trend across distinct $(a,C)$ families is captured by a single fit (cyan dashed line). Remarkably, the same fitting form successfully described strange stars in previous work [Lugones2025universality] based on the flavor-blind QMDD and vector MIT bag models. Its effectiveness here, for stars with different EOS ingredients and internal compositions, underscores its high degree of universality.
@@
It is worth emphasizing that the $\bar{I}$$\mathscr{C}$ relation for self-bound quark stars found here differs markedly from the purely hadronic relation of Breu and Rezzolla [Breu:2016ufb]. Consequently, a simultaneous measurement of $M$, $R$, and $I$ would place a source at distinct loci in the $\bar{I}$$\mathscr{C}$ plane, offering a clean diagnostic to discriminate a self-bound quark star from a conventional hadronic star.
VI.2 Tidal deformability versus compactness
Figure 11(a) displays an approximately universal relation between the dimensionless tidal deformability $\Lambda$ and the compactness $\mathscr{C}=M/R$ for the self-bound quark-star models considered. In close analogy with the $\bar{I}$$\mathscr{C}$ relation, the dependence on the excluded-volume parameter is negligible: for a fixed $(a,C)$, sequences with $\kappa=0,\,0.3,$ and $0.6$ are practically indistinguishable and collapse onto a single curve. Moreover, curves from distinct $(a,C)$ families cluster tightly around the strange-star fit (cyan dashed) obtained in Ref. [Lugones2025universality] using the flavor-blind parametrization. The fidelity of this universal behavior is quantified in Fig. 11(b), which shows the relative deviation from the fit.
For reference, Fig. 11(a) also includes a representative hadronic fit (crimson dashed). The two relations remain clearly distinct over the full compactness range: at fixed $\mathscr{C}$, hadronic stars exhibit systematically larger $\Lambda$ than self-bound quark stars. This offset has a clear physical origin. Self-bound objects possess a finite surface density and more homogeneous interiors, which yield smaller Love numbers $k_{2}$; because $\Lambda\propto k_{2}\,\mathscr{C}^{-5}$, the reduced $k_{2}$ directly translates into a smaller tidal deformability.
In summary: (i) the $\Lambda$$\mathscr{C}$ relation is effectively insensitive to $\kappa$; (ii) the strange-star fit from Ref. [Lugones2025universality]—previously shown to describe vMIT + QMDD flavor-blind data—also captures models with internal $ud\!\to\!uds$ transitions; and (iii) the quark-star and hadronic relations are well separated, offering an observational discriminator across the entire compactness range explored.
VI.3 Gravitational versus baryonic compactness
Figure 12: Gravitational compactness $\mathscr{C}$ versus baryonic compactness $\mathscr{C}_{B}$ for stellar sequences labeled by the parameter $a$ (colors as in the legend). Panel (a): solid curves show the models, while the cyan dashed line is the strange–quark–star fit [Lugones2025universality]. Panel (b): relative error with respect to the fit, $|1-\mathscr{C}/\mathscr{C}^{\mathrm{fit}}|$.

### v2

@@ IV NUMERICAL RESULTS FOR THE EOS
Before presenting the detailed numerical results, it is useful to clarify the physical origin of the $ud\!\to\!uds$ transition in the present framework. In the flavor-dependent QMDD formulation, each dynamical quark mass tracks the number density of its own flavor, $M_{i}=M_{i}(n_{i})$, rather than a common baryon-density scale as in flavor-blind variants. This weakens the correlation among the density evolutions of $M_{u}$, $M_{d}$, and $M_{s}$, so that strangeness can remain strongly suppressed at low densities (large $M_{s}$) while $u$ and $d$ quarks are not suppressed to the same extent. Combined with our excluded-volume prescription, this mechanism naturally generates a stability window in which self-bound two-flavor matter exists at $p\!\to\!0$ even though three-flavor matter is not yet favored. As the density increases, $M_{s}$ decreases and udsuds matter eventually becomes thermodynamically preferred; the conversion occurs when the Gibbs free energies cross, producing a genuine first-order transition with a finite density (energy-density) jump. Importantly, the transition is not imposed by hand, but emerges dynamically from the density dependence of the quark masses.
@@
Table 2: First-order $ud\!\to\!uds$ transition parameters for the flavor-dependent QMDD EOS. For each $(a,C,\kappa)$ we list the transition pressure $p_{\mathrm{tr}}$, the energy densities just below/above the transition ($\epsilon_{-}$, $\epsilon_{+}$), and the jump $\Delta\epsilon\equiv\epsilon_{+}-\epsilon_{-}$. The last column reports the Seidov onset-instability diagnostic $\gamma$ defined in Eq. (7); values $\gamma<1$ indicate that Seidov’s threshold is not reached at the transition onset. We emphasize that the definitive stability assessment is obtained from the TOV equilibrium sequences (Sec. V), and we use $\gamma$ only as an auxiliary check.
$a$ $C$ $\kappa$ $p_{\mathrm{tr}}$ (MeV fm-3) $\epsilon_{+}$ (MeV fm-3) $\epsilon_{-}$ (MeV fm-3) $\Delta\epsilon$ (MeV fm-3) $\gamma$
@@
Seidov [30] derived a useful onset-instability diagnostic for first-order phase transitions in general relativity: a spherically symmetric configuration becomes unstable at the transition onset if the energy-density discontinuity across the interface exceeds the critical threshold
@@
where $p_{\mathrm{tr}}$ is the transition pressure and $\epsilon_{-}$ ($\epsilon_{+}$) is the energy density just below (above) the transition. For convenience, Table 2 lists $p_{\mathrm{tr}}$, $\epsilon_{-}$, and $\epsilon_{+}$ for the benchmark EOSs considered here and reports in the last column the corresponding diagnostic ratio
$\gamma\equiv\frac{\epsilon_{+}/\epsilon_{-}}{\tfrac{3}{2}\bigl(1+p_{\mathrm{tr}}/\epsilon_{-}\bigr)}\,.$ (7)
We find $\gamma<1$ for all benchmark transitions shown (typically 0.850.850.970.97), indicating that Seidov’s onset-instability threshold is not reached in these cases. We stress, however, that this criterion is only an auxiliary check: the definitive stability assessment is obtained by constructing the full TOV sequences and applying the turning-point condition. We confirm in Sec. V that the sequences containing a $ud\!\to\!uds$ interface exhibit dynamically stable hybrid branches.
Table 2 further shows that, for fixed $(a,C)$, the individual transition quantities $p_{\mathrm{tr}}$, $\epsilon_{-}$, and $\epsilon_{+}$ vary appreciably with $\kappa$, whereas the Seidov diagnostic $\gamma$ is essentially insensitive to $\kappa$ within each family. Finally, we have verified that $\gamma$ remains below unity throughout the causal self-bound region of the $(a,C)$ parameter space explored in Fig. 1; since this diagnostic is not central to the main goals of this work, we do not include an exhaustive scan here, but note it for completeness.
@@ V Global properties of Hybrid self-bound quark stars
Figure 5: (a) Schematic mass–radius diagram of self-bound stars. The magenta curve shows self-bound strange-quark-matter configurations, the blue curve shows self-bound udud stars, and the bicolor (magenta–blue) curve displays the sequence of self-bound hybrid stars analyzed in this work. (b) Representation of the internal composition of the models on the MMRR curves in panel (a), indicating regions of udsuds matter (magenta) and $ud$ matter (blue).
@@
Figure 6: Mass–radius relations for quark-star sequences computed with the EOS sets of Table 1. Line styles encode the excluded-volume strength $\kappa$ (solid, dashed, dotted), following the convention of Fig. 2. Colors indicate composition: magenta—pure udsuds (SQM); blue—pure udud; bicolor—self-bound hybrids with a udsuds core surrounded by a $ud$ mantle. (a) Parameters in the SQM domain yield purely three-flavor sequences. (b)–(d) EOSs featuring a first-order $ud\!\to\!uds$ transition placed at low, intermediate, and high densities, respectively; the kink marks $p_{c}=p_{\mathrm{tr}}$. Increasing $\kappa$ stiffens the EOS and raises $M_{\max}$. Shaded contours denote observational constraints from PSR J0030+0451 [28, 23], HESS J1731-347 [6], and PSR J0740+6620 [29, 24].
@@
Figure 6 confronts the mass–radius sequences with current astrophysical constraints. A clear dependence on the excluded-volume strength is evident. Models with $\kappa=0$ are too soft: their MMRR curves fail to reach $2\,M_{\odot}$, underscoring the need for repulsive interactions in the quark sector. At the other extreme, $\kappa=0.6$ produces very stiff EOSs. Although these sequences comfortably attain large maximum masses, they generally overpredict radii; in panels 6(b)–6(d) the curves lie on or above the upper edge of the PSR J0030+0451 and PSR J0740+6620 constraints. An exception occurs in panel 6(a), where the pure SQM models with $\kappa=0.6$ are compatible with all bands, a marked improvement over $\kappa=0$ and $0.3$. For the self-bound hybrid cases, the intermediate choice $\kappa=0.3$ provides the best overall agreement, balancing the $2\,M_{\odot}$ requirement against radius constraints. Across panels 6(b)–6(d), the predicted $M_{\max}$ varies with the transition density: it is lower for both low and high transition densities and peaks for the intermediate case in panel 6(c).
Figure 7: Dimensionless tidal deformability $\Lambda$ as a function of gravitational mass $M$ for the same EOS sets used in Fig. 6. Line styles encode the excluded-volume strength $\kappa$ (solid, dashed, dotted), following the convention of Fig. 2. (a) Pure strange–quark–star sequences; (b)–(d) self-bound hybrid-star sequences with the $ud\!\to\!uds$ transition placed at low, intermediate, and high densities, respectively.
@@
Figure 7 displays the (dimensionless) tidal deformability $\Lambda$, which quantifies the induced quadrupole moment per unit external tidal field and is a key finite-size observable of compact stars in inspiralling binaries. We compute $\Lambda(M)$ for the same four EOS sets as in the preceding figures, using the formalism summarized in Appendix B.2. Two robust trends emerge. First, for any given EOS, $\Lambda$ decreases monotonically with $M$ as stars become more compact. Second, at fixed $M$, a larger excluded-volume strength (larger $\kappa$) yields larger $\Lambda$, reflecting the stiffening of the EOS and the associated increase in stellar radius. Quantitatively, one may write $\Lambda=\tfrac{2}{3}\,k_{2}\,\mathscr{C}^{-5}\simeq\tfrac{2}{3}\,k_{2}\,(R/M)^{5}$, so that even modest radius variations at fixed mass can translate into sizeable changes in $\Lambda$ through its steep compactness dependence.
The two classes of models exhibit distinct behaviors. For pure strange-quark stars [Fig. 7(a)], $\Lambda$ remains comparatively small across the mass range, reflecting their higher compactness. By contrast, self-bound hybrid stars [Figs. 7(b)–7(d)] typically produce larger $\Lambda$ at a fixed $M$, with the detailed level set by the transition density and by the repulsive excluded-volume correction: increasing $\kappa$ systematically shifts the curves upward, whereas varying the transition density alters the location and extent of the hybrid branch. For configurations featuring a sharp first–order interface and/or a finite surface density (self-bound stars), the tidal perturbation variable $y(r)$ is matched using the junction prescriptions detailed in Appendix B.2, Eqs. (42) and (43), before evaluating $k_{2}$ and $\Lambda$.
It is instructive to compare our $\Lambda(M)$ curves with the constraint reported in Ref. [1]. In the low-spin scenario and assuming a uniform prior on the effective tidal parameter $\tilde{\Lambda}$, that analysis yields a $90\%$ upper limit $\tilde{\Lambda}\leq 800$; expanding $\Lambda(m)$ linearly about $1.4\,M_{\odot}$ further gives $\Lambda(1.4\,M_{\odot})\leq 800$. We adopt this single-star threshold as an approximate observational reference and indicate it by the gray markers in each panel of Fig. 7. For the pure strange-quark-star sequences [Fig. 7(a)], the models with $\kappa=0$ and $\kappa=0.3$ predict $\Lambda(1.4\,M_{\odot})$ comfortably below the observational ceiling, reflecting the high compactness characteristic of self-bound udsuds stars. The stiffest case, $\kappa=0.6$, yields a somewhat larger $\Lambda$ but still remains compatible with the bound. In the self-bound hybrid case, the constraint is considerably more restrictive. In panels 7(b) and 7(c), only the softest parametrization ($\kappa=0$) satisfies $\Lambda(1.4\,M_{\odot})\lesssim 800$, and it does so only marginally, with the curve lying just below the observational threshold. Both $\kappa=0.3$ and $\kappa=0.6$ exceed the bound in these panels, the latter by a wide margin. The situation improves markedly for the high-transition-density set [panel 7(d)], where the later onset of the udsuds core keeps the stellar radius smaller at $1.4\,M_{\odot}$: here both $\kappa=0$ and $\kappa=0.3$ lie well below $800$, while only the $\kappa=0.6$ models violate the constraint. Physically, these trends reflect the steep compactness dependence of the tidal deformability, $\Lambda\propto\mathscr{C}^{-5}$: a stiffer EOS inflates the stellar radius, and even a modest increase is amplified into a large enhancement of $\Lambda$. Notably, in panels 7(b)–7(c) the $\kappa=0$ models are the only ones compatible with the GW170817 tidal bound, yet these same parametrizations were already ruled out in the mass–radius analysis of Fig. 6 for failing to support $2\,M_{\odot}$ stars. This combined tension disfavors the low- and intermediate-transition-density EOS sets for the self-bound hybrid scenario. In contrast, the high-transition-density set [panel 7(d)] with $\kappa=0.3$ simultaneously satisfies the tidal constraint and the $2\,M_{\odot}$ requirement, confirming the conclusions drawn from Fig. 6 and identifying this region of parameter space as the most observationally viable for self-bound hybrid configurations.
@@ VI Universal relations
Figure 10: (a) Dimensionless moment of inertia $I/M_{G}^{3}$ vs. compactness $M_{G}/R$ for sequences with different values of $a$. Only EOSs with a maximum mass $2\leq M_{\max}/M_{\odot}\leq 3$ are included. The dashed lines are the global fits for strange quark stars [19] and hadronic stars [3]. (b) Relative error, $\bigl|1-(I/M_{G}^{3})_{\rm num}/(I/M_{G}^{3})_{\rm fit}\bigr|$, with respect to the strange quark star fit.
@@ VI.1 Dimensionless moment of inertia
Figure 10 shows the dimensionless moment of inertia, $\bar{I}\equiv I/M^{3}$, as a function of the compactness $\mathscr{C}\equiv M/R$. Despite the strong $\kappa$-dependence seen in individual MRM\!-\!R and MIM\!-\!I relations, its imprint is nearly washed out in the $\bar{I}$$\mathscr{C}$ plane: for fixed $(a,C)$, the three $\kappa$ tracks collapse to an excellent approximation onto a single curve. Moreover, the global trend across distinct $(a,C)$ families is captured by a single fit (cyan dashed line). Remarkably, the same fitting form successfully described strange stars in previous work [19] based on the flavor-blind QMDD and vector MIT bag models. Its effectiveness here, for stars with different EOS ingredients and internal compositions, underscores its high degree of universality.
@@
It is worth emphasizing that the $\bar{I}$$\mathscr{C}$ relation for self-bound quark stars found here differs markedly from the purely hadronic relation of Breu and Rezzolla [3]. Consequently, a simultaneous measurement of $M$, $R$, and $I$ would place a source at distinct loci in the $\bar{I}$$\mathscr{C}$ plane, offering a clean diagnostic to discriminate a self-bound quark star from a conventional hadronic star.
VI.2 Gravitational versus baryonic compactness
Figure 11: Gravitational compactness $\mathscr{C}$ versus baryonic compactness $\mathscr{C}_{B}$ for stellar sequences labeled by the parameter $a$ (colors as in the legend). Panel (a): solid curves show the models, while the cyan dashed line is the strange–quark–star fit [19]. Panel (b): relative error with respect to the fit, $|1-\mathscr{C}/\mathscr{C}^{\mathrm{fit}}|$.